# On Jiang’s wavefront sets conjecture for representations in local Arthur packets

### 13 Mar 2026 v1

@@ 2. Characterization of local Arthur packets
At the end of this section, we give the proof for Theorem 1.1 and give a generalization of Conjecture 1.2.
@@ 4. Dimension identities for nilpotent orbits
where $f\in\widetilde{{\mathcal{H}}}(G_{n})$, $f^{\prime}\in\widetilde{{\mathcal{S}}}(G^{\prime})$ is the transfer of $f$ to $G^{\prime}$ with the assumption that $f^{\prime}=f^{\prime 1}\otimes f^{\prime 2}$, $f^{\prime i}\in\widetilde{{\mathcal{S}}}(G^{\prime}_{i})$, $i=1,2$. Let
$\widetilde{f}^{1}\in\begin{cases}\widetilde{{\mathcal{H}}}(2n_{1}+1)&\text{ when }\mathrm{G}_{n}={\mathrm{Sp}}_{2n},\\ \widetilde{{\mathcal{H}}}(2n_{1})&\text{ when }\mathrm{G}_{n}={\mathrm{SO}}_{2n+1},{\mathrm{SO}}_{2n}^{\alpha},\end{cases}$
and $\widetilde{f}^{2}\in\widetilde{{\mathcal{H}}}(2n_{2})$, transferring to $f^{\prime 1}$ and $f^{\prime 2}$, via surjective maps in [Art13, Corollary 2.1.2]
$\iota_{G_{k}}:\widetilde{{\mathcal{H}}}(N)\rightarrow\widetilde{{\mathcal{S}}}(G_{k}),k=n_{1},n_{2},$
@@ 4.1. Character expansions
In this section, we take the character expansions of both sides of (4.1). For the left hand side, as in [HC78] for connected groups, we take the character expansions of characters at the center points $\pm Id$. For the right hand side, as in [Clo87] for disconnected groups, we take the character expansion of the twisted characters at $\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}$ (see (1.3)), $i=1,2$, or at $\theta_{\widehat{{\mathrm{G}}}_{1}^{\prime}}$ (see (1.3)) when ${\mathrm{G}}_{1}^{\prime}={{\mathrm{Sp}}}_{2n_{1}}$. We discuss the details by cases.
Case $\mathrm{G}_{n}={\mathrm{Sp}}_{2n}$. Recall that ${\mathrm{G}}_{1}^{\prime}={\mathrm{Sp}}_{2n_{1}}$, ${\mathrm{G}}_{2}^{\prime}={\mathrm{SO}}_{2n_{2}}^{\beta}$, $\widehat{{\mathrm{G}}}_{1}^{\prime}(\mathbb{C})={\mathrm{SO}}_{2n_{1}+1}({\mathbb{C}})$, $\widehat{{\mathrm{G}}}_{2}^{\prime}(\mathbb{C})={\mathrm{SO}}_{2n_{2}}({\mathbb{C}})$. Also recall that
$\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}=s_{\widehat{\mathrm{G}}_{1}^{\prime}}\rtimes\widetilde{\theta}({2n_{1}+1})\in\widetilde{{\mathrm{GL}}}_{2n_{1}+1}(F),$
where
$s_{\widehat{\mathrm{G}}_{1}^{\prime}}=\begin{pmatrix}I_{n_{1}}&&\\ &1&\\ &&-I_{n_{1}}\end{pmatrix}$.
The norm of $s_{\widehat{\mathrm{G}}_{1}^{\prime}}$ is
$s_{\widehat{\mathrm{G}}_{1}^{\prime}}\cdot\widetilde{\theta}(2n_{1}+1)(s_{\widehat{\mathrm{G}}_{1}^{\prime}})=\begin{pmatrix}-I_{n_{1}}&&\\ &1&\\ &&-I_{n_{1}}\end{pmatrix}$.
$\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}=s_{\widehat{\mathrm{G}}_{2}^{\prime}}\rtimes\widetilde{\theta}({2n_{2}})\in\widetilde{{\mathrm{GL}}}_{2n_{2}}(F),$
where
$s_{\widehat{\mathrm{G}}_{2}^{\prime}}=\begin{pmatrix}I_{n_{2}}&\\ &-I_{n_{2}}\end{pmatrix}$.
The norm of $s_{\widehat{\mathrm{G}}_{2}^{\prime}}$ is
$s_{\widehat{\mathrm{G}}_{2}^{\prime}}\cdot\widetilde{\theta}(2n_{2})(s_{\widehat{\mathrm{G}}_{2}^{\prime}})=\begin{pmatrix}-I_{n_{2}}&\\ &-I_{n_{2}}\end{pmatrix}$.
Therefore, we can take the character expansions of characters on the left hand side of (4.1) at $I_{2n}$ and take the character expansion of the twisted characters ${\mathrm{tr}}(\widetilde{\pi}_{\psi^{i}}(\widetilde{f}^{i}))$ on the right hand side of (4.1) at $\theta_{\widehat{{\mathrm{G}}}_{i}^{\prime}}$ (see (1.3)), $i=1,2$. Note that taking the character expansion of $f_{G_{n}}(\pi)=tr(\pi(f))$ at $I_{2n}$ is equivalent to taking the character expansion of $\omega_{\pi}(-I_{2n})tr(\pi(R(-I_{2n})f))$ at $I_{2n}$, where $R(-I_{2n})f$ is the right translation of $f$ by $I_{2n}$.
Notice that the connected component of the stabilizer of $\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}$ in $\mathrm{GL}_{2n_{1}+1}$ is ${\mathrm{G}}_{1}^{\prime}\times{\mathrm{SO}}_{1}$ and the connected component of the stabilizer of $\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}$ in $\mathrm{GL}_{2n_{2}}$ is $\widehat{\mathrm{G}}_{2}^{\prime}$.
Let $c_{{\mathrm{G}}_{n}}=-I_{2n}$. After taking the character expansions as above we have the following equality:
@@
(4.2) $\begin{split}&\,\sum_{\pi\in\widetilde{\Pi}_{\psi}}\langle 1,\pi\rangle\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}\omega_{\pi}(c_{{\mathrm{G}}_{n}})c_{{\mathcal{O}}}(\pi)\hat{\mu}_{\mathcal{O}}(R(c_{{\mathrm{G}}_{n}})f)\\ =&\,\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{{\mathfrak{g}}_{1}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}})\right)\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{2}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}})\right),\end{split}$
@@
where ${\mathcal{N}}_{\mathfrak{g}}$ denotes the set of $F$-rational nilpotent orbits in a Lie algebra $\mathfrak{g}$, $\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}}^{i}$ is the Harish-Chandra descent of $\widetilde{f}^{i}$ (see for example [Kon02, Section 3.1]), $i=1,2$. $R(c_{{\mathrm{G}}_{n}})f$ denotes the right translation of $f$ by $c_{{\mathrm{G}}_{n}}$. Equation (4.2) can be rewritten as follows:
@@
(4.3) $\begin{split}&\,\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{\mathcal{O}}(R(c_{{\mathrm{G}}_{n}})f)\\ =&\,\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{{\mathfrak{g}}_{1}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}})\right)\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{2}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}})\right),\end{split}$
where $c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})=\sum_{\pi\in\widetilde{\Pi}_{\psi}}\langle 1,\pi\rangle\omega_{\pi}(c_{{\mathrm{G}}_{n}})c_{{\mathcal{O}}}(\pi)$.
Note that by Conjecture 1.2 and Remark 1.4(4), the maximal nilpotent orbits occurring on the right hand side of (4.3) are those corresponding to the partitions
$(([\prod_{i\in I}b_{i}^{a_{i}}]^{t})^{-})_{{\mathrm{G}}_{1}^{\prime}}=(([\prod_{i\in I}b_{i}^{a_{i}}]^{-})_{{\mathrm{G}}_{1}^{\prime}})^{t}=\eta_{{\hat{\mathfrak{g}}_{1}^{\prime},\mathfrak{g}_{1}^{\prime}}}(\underline{p}(\psi^{1}))$
and $([\prod_{j\in J}b_{j}^{a_{j}}]^{t})_{\widehat{\mathrm{G}}_{2}^{\prime}}$.
Case $\mathrm{G}_{n}={\mathrm{SO}}_{2n+1}$. Recall that ${\mathrm{G}}_{1}^{\prime}={\mathrm{SO}}_{2n_{1}+1}$, ${\mathrm{G}}_{2}^{\prime}={\mathrm{SO}}_{2n_{2}+1}$, $\widehat{{\mathrm{G}}}_{1}^{\prime}(\mathbb{C})={\mathrm{Sp}}_{2n_{1}}({\mathbb{C}})$, $\widehat{{\mathrm{G}}}_{2}^{\prime}(\mathbb{C})={\mathrm{Sp}}_{2n_{2}}({\mathbb{C}})$. Also recall that
$\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}=s_{\widehat{\mathrm{G}}_{i}^{\prime}}\rtimes\widetilde{\theta}({2n_{i}})\in\widetilde{{\mathrm{GL}}}_{2n_{i}}(F),$
where $s_{\widehat{\mathrm{G}}_{i}^{\prime}}=I_{2n_{i}}$, $i=1,2$. Therefore, we can take the character expansions of characters on the left hand side of (4.1) at $I_{2n+1}$ and take the character expansion of the twisted characters ${\mathrm{tr}}(\widetilde{\pi}_{\psi^{i}}(\widetilde{f}^{i}))$ on the right hand side of (4.1) at $\theta_{\widehat{{\mathrm{G}}}_{i}^{\prime}}$ (see (1.3)), $i=1,2$.
Notice that the connected component of the stabilizer of $\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}$ in $\mathrm{GL}_{2n_{i}}$ is $\widehat{\mathrm{G}}_{i}^{\prime}$. Let $c_{{\mathrm{G}}_{n}}=I_{2n+1}$. After taking the character expansions as above, we have the following equality:
(4.4) $\begin{split}&\,\sum_{\pi\in\widetilde{\Pi}_{\psi}}\langle 1,\pi\rangle\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}\omega_{\pi}(c_{{\mathrm{G}}_{n}})c_{{\mathcal{O}}}(\pi)\hat{\mu}_{\mathcal{O}}(R(c_{{\mathrm{G}}_{n}})f)\\ =&\,\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{1}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}})\right)\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{2}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}})\right).\end{split}$
Equation (4.4) can be rewritten as follows:
(4.5) $\begin{split}&\,\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{\mathcal{O}}(R(c_{{\mathrm{G}}_{n}})f)\\ =&\,\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{1}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}})\right)\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{2}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}})\right),\end{split}$
where $c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})=\sum_{\pi\in\widetilde{\Pi}_{\psi}}\langle 1,\pi\rangle\omega_{\pi}(c_{{\mathrm{G}}_{n}})c_{{\mathcal{O}}}(\pi)$.
Note that by Conjecture 1.2, the maximal nilpotent orbits occurring on the right hand side of (4.5) are those corresponding to the partitions $([\prod_{i\in I}b_{i}^{a_{i}}]^{t})_{\widehat{\mathrm{G}}_{1}^{\prime}}$ and $([\prod_{j\in J}b_{j}^{a_{j}}]^{t})_{\widehat{\mathrm{G}}_{2}^{\prime}}$.
Case $\mathrm{G}_{n}={\mathrm{SO}}_{2n}^{\alpha}$. Recall that ${\mathrm{G}}_{1}^{\prime}={\mathrm{SO}}_{2n_{1}}^{\gamma_{1}}$, ${\mathrm{G}}_{2}^{\prime}={\mathrm{SO}}_{2n_{2}}^{\gamma_{2}}$, $\widehat{{\mathrm{G}}}_{1}^{\prime}(\mathbb{C})={\mathrm{SO}}_{2n_{1}}({\mathbb{C}})$, $\widehat{{\mathrm{G}}}_{2}^{\prime}(\mathbb{C})={\mathrm{SO}}_{2n_{2}}({\mathbb{C}})$. Also recall that
$\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}=s_{\widehat{\mathrm{G}}_{i}^{\prime}}\rtimes\widetilde{\theta}({2n_{i}})\in\widetilde{{\mathrm{GL}}}_{2n_{i}}(F),$
where
$s_{\widehat{\mathrm{G}}_{i}^{\prime}}=\begin{pmatrix}I_{n_{i}}&\\ &-I_{n_{i}}\end{pmatrix},i=1,2.$
The norm of $s_{\widehat{\mathrm{G}}_{i}^{\prime}}$ is
$s_{\widehat{\mathrm{G}}_{i}^{\prime}}\cdot\widetilde{\theta}(2n_{i})(s_{\widehat{\mathrm{G}}_{i}^{\prime}})=\begin{pmatrix}-I_{n_{i}}&\\ &-I_{n_{i}}\end{pmatrix}$.
Therefore, we can take the character expansions of characters on the left hand side of (4.1) at $I_{2n}$ and take the character expansion of the twisted characters ${\mathrm{tr}}(\widetilde{\pi}_{\psi^{i}}(\widetilde{f}^{i}))$ on the left hand side of (4.1) at $\theta_{\widehat{{\mathrm{G}}}_{i}^{\prime}}$ (see (1.3)), $i=1,2$. Again, taking the character expansion of $f_{G_{n}}(\pi)=tr(\pi(f))$ at $I_{2n}$ is equivalent as taking the character expansion of $\omega_{\pi}(-I_{2n})tr(\pi(R(-I_{2n})f))$ at $I_{2n}$.
Notice that the connected component of the stabilizer of $\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}$ in $\mathrm{GL}_{2n_{i}}$ is $\widehat{\mathrm{G}}_{i}^{\prime}$. Let $c_{{\mathrm{G}}_{n}}=-I_{2n}$. After taking the character expansions as above we have the following equality:
(4.6) $\begin{split}&\,\sum_{\pi\in\widetilde{\Pi}_{\psi}}\langle 1,\pi\rangle\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}\omega_{\pi}(c_{{\mathrm{G}}_{n}})c_{{\mathcal{O}}}(\pi)\hat{\mu}_{\mathcal{O}}(R(c_{{\mathrm{G}}_{n}})f)\\ =&\,\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{1}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}})\right)\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{2}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}})\right).\end{split}$
Equation (4.6) can be rewritten as follows:
(4.7) $\begin{split}&\,\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{\mathcal{O}}(R(c_{{\mathrm{G}}_{n}})f)\\ =&\,\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{1}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}})\right)\left(\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\hat{\mathfrak{g}}_{2}^{\prime}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}})\right),\end{split}$
where $c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})=\sum_{\pi\in\widetilde{\Pi}_{\psi}}\langle 1,\pi\rangle\omega_{\pi}(c_{{\mathrm{G}}_{n}})c_{{\mathcal{O}}}(\pi)$.
Note that by Conjecture 1.2, the maximal nilpotent orbits occurring on the right hand side of (4.7) are those corresponding to the partitions $([\prod_{i\in I}b_{i}^{a_{i}}]^{t})_{\widehat{\mathrm{G}}_{1}^{\prime}}$ and $([\prod_{j\in J}b_{j}^{a_{j}}]^{t})_{\widehat{\mathrm{G}}_{2}^{\prime}}$.

### 06 May 2026 v2

@@ 2. Characterization of local Arthur packets
At the end of this section, we give the proof for Theorem 1.1.
@@ 4. Dimension identities for nilpotent orbits
where $f\in\widetilde{{\mathcal{H}}}(G_{n})$, $f^{\prime}\in\widetilde{{\mathcal{S}}}(G^{\prime})$ is the transfer of $f$ to $G^{\prime}$ with the assumption that $f^{\prime}=f^{\prime 1}\otimes f^{\prime 2}$, $f^{\prime i}\in\widetilde{{\mathcal{S}}}(G^{\prime}_{i})$, $i=1,2$. Let $N_{1}=\begin{cases}2n_{1}+1&\text{ when }\mathrm{G}_{n}={\mathrm{Sp}}_{2n},\\ 2n_{1}&\text{ when }\mathrm{G}_{n}={\mathrm{SO}}_{2n+1},{\mathrm{SO}}_{2n}^{\alpha},\end{cases}$ and $N_{2}=2n_{2}$. Let $\widetilde{f}^{1}\in\widetilde{{\mathcal{H}}}(N_{1})$ and $\widetilde{f}^{2}\in\widetilde{{\mathcal{H}}}(N_{2})$, transferring to $f^{\prime 1}$ and $f^{\prime 2}$, via surjective maps in [Art13, Corollary 2.1.2]
$\iota_{G^{\prime}_{i}}:\widetilde{{\mathcal{H}}}(N_{i})\rightarrow\widetilde{{\mathcal{S}}}(G^{\prime}_{i}),i=1,2,$
@@ 4.1. Character expansions
In this section, we take the character expansions of both sides of (4.1). The left hand side is expanded at the identity, and the right hand side is expanded at the twisted points $\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}=1\rtimes\widetilde{\theta}(N_{i})$ (see (1.3)), $i=1,2$. Then we have the following equality:
@@
(4.2) $\begin{split}&\,\sum_{\pi\in\widetilde{\Pi}_{\psi}}\langle 1,\pi\rangle\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}c_{{\mathcal{O}}}(\pi)\hat{\mu}_{\mathcal{O}}(f)\\ =&\,\left(\sum{{}_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g_{1}^{\prime}}^{\theta}}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}})\right)\left(\sum{{}_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g_{2}^{\prime}}^{\theta}}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}})\right),\end{split}$
@@
where ${\mathcal{N}}_{\mathfrak{g_{i}^{\prime}}^{\theta}}$ denotes the set of $F$-rational nilpotent orbits in the Lie algebra $\mathfrak{g_{i}^{\prime}}^{\theta}$, $\widetilde{f}_{\theta_{\widehat{\mathrm{G}}_{i}^{\prime}}}^{i}$ is the Harish-Chandra descent of $\widetilde{f}^{i}$ (see for example [Kon02, Section 3.1]), $i=1,2$. Equation (4.2) can be rewritten as follows:
@@
(4.3) $\begin{split}&\,\sum_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g}_{n}}}c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})\hat{\mu}_{\mathcal{O}}(f)\\ =&\,\left(\sum{{}_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g_{1}^{\prime}}^{\theta}}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{1}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{1}_{\theta_{\widehat{\mathrm{G}}_{1}^{\prime}}}})\right)\left(\sum{{}_{{\mathcal{O}}\in{\mathcal{N}}_{\mathfrak{g_{2}^{\prime}}^{\theta}}}}c_{{\mathcal{O}}}(\widetilde{\pi}_{\psi^{2}})\hat{\mu}_{\mathcal{O}}({{}\widetilde{f}^{2}_{\theta_{\widehat{\mathrm{G}}_{2}^{\prime}}}})\right),\end{split}$
where $c_{{\mathcal{O}}}(\widetilde{\Pi}_{\psi})=\sum_{\pi\in\widetilde{\Pi}_{\psi}}c_{{\mathcal{O}}}(\pi)$. Note that ${{\mathrm{G}}^{\prime}_{1}}^{\theta}=\begin{cases}{\mathrm{SO}}_{2n_{1}+1}&\text{ when }\mathrm{G}_{n}={\mathrm{Sp}}_{2n},\\ {\mathrm{Sp}}_{2n_{1}}&\text{ when }\mathrm{G}_{n}={\mathrm{SO}}_{2n+1},{\mathrm{SO}}_{2n}^{\alpha},\end{cases}$, ${{\mathrm{G}}^{\prime}_{2}}^{\theta}={\mathrm{Sp}}_{2n_{2}}$, and $\mathfrak{g_{i}^{\prime}}^{\theta}$ is the Lie algebra of ${{\mathrm{G}}^{\prime}_{i}}^{\theta}$.